OpenStudy (anonymous):

9+ |-4|= I don't get it.

OpenStudy (anonymous):

13

OpenStudy (anonymous):

whatever is inside the absolute value bars becomes positive, always

OpenStudy (anonymous):

ex ......2+l-5l = 2+5 = 7

OpenStudy (anonymous):

if u have absolute value the number will always be positive so its gonna be 13

OpenStudy (anonymous):

the two vertical lines mean absolute value of whats inside or sometimes called 'mod -4'